Given: [a/(b+c)] + [b(a+c)] + [c/(a+b)] = 1 Then What is the Value of : [a 2 /(b+c)] + [b 2 (a+c)] + [c 2 /(a+b)] ?

Given:
 
[a/(b+c)] + [b(a+c)] + [c/(a+b)] = 1
 
Then What is the Value of :
[a2/(b+c)] + [b2(a+c)] + [c2/(a+b)] ?
 

Grade:11

3 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the response to your question below
I think there is some data missing or some mistake in your question.
whether it is b(a+c) or b/a+c also it is b2(a+c) or b2/a+c ?
please rechek the question and post it again so that i can provide you with a meaningful answer.
Hari Eeshwar
9 years ago
Sir,
Thankyou for the response so that i could post the accurate question as follows:
[a/(b+c)] + [b/(a+c)] + [c/(a+b)] = 1
 
Then What is the Value of :
[a2/(b+c)] + [b2/(a+c)] + [c2/(a+b)] =?
mycroft holmes
9 years ago
We have (a+b+c)[ (a/b+c) + (b/c+a) + (c/a+b) ] = a^2/(b+c)+b^2/(c+a)+c^2/(a+b) + (a+b+c)
 
Hence if (a/b+c) + (b/c+a) + (c/a+b) =1, the equation becomes
a+b+c = a^2/(b+c)+b^2/(c+a)+c^2/(a+b) + (a+b+c) 
 
so that a^2/(b+c)+b^2/(c+a)+c^2/(a+b) = 0